The product of four greater than an certain negative integer and sixteen less than four times the integers is equal to 36.Determine the integer ?
Answer by stanbon(75887)

The product of four greater than an certain negative integer and sixteen less than four times the integers is equal to 36.Determine the integer ?
-----------------------
(x+4)(4x-16) = 36
(x+4)(x-4) = 9
x^2-16 = 9
x^2 = 25
x = 5 or x = -5
======================
Cheers,
Stan H.